OpenStudy (anonymous):

if csc x= -5/4 and 180

jimthompson5910 (jim_thompson5910):

First find x in 

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

I guess I should have used another variable than x, but tan(x) = opp/adj tan(x) = -4/5

OpenStudy (anonymous):

isnt 5 the hypotenuse?

OpenStudy (anonymous):

would it be -4/-3 since its in the 3rd quadrant making it just 4/3? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

oh right, yeah it's positive, my bad

jimthompson5910 (jim_thompson5910):

and it should be tan(x) = 4/3

jimthompson5910 (jim_thompson5910):

so you are 100% correct

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

np